### v1

# Synthesizing Moving People with 3D Control

Abstract
In this paper, we present a diffusion model-based framework for animating people from a single image for a given target 3D motion sequence. Our approach has two core components: a) learning priors about invisible parts of the human body and clothing, and b) rendering novel body poses with proper clothing and texture. For the first part, we learn an in-filling diffusion model to hallucinate unseen parts of a person given a single image. We train this model on texture map space, which makes it more sample-efficient since it is invariant to pose and viewpoint. Second, we develop a diffusion-based rendering pipeline, which is controlled by 3D human poses. This produces realistic renderings of novel poses of the person, including clothing, hair, and plausible in-filling of unseen regions. This disentangled approach allows our method to generate a sequence of images that are faithful to the target motion in the 3D pose and, to the input image in terms of visual similarity. In addition to that, the 3D control allows various synthetic camera trajectories to render a person. Our experiments show that our method is resilient in generating prolonged motions and varied challenging and complex poses compared to prior methods. Please check our website for more details: 3DHM.github.io.
Figure 1: The Imitation Game: Given a video of a person "The Actor", we want to transfer their motion to a new person "The Imitator". In this figure, the first row shows a sequence of frames of the actor (Michelle Kwan), doing her Olympics ’98 performance. The inset row shows the 3D poses extracted from this video. Now, given any single image of a new person The Imitator, our model can synthesize new renderings of the imitator, to copy the actions of the actor in 3D.
1 Introduction
Given a random photo of a person, can we accurately animate that person to imitate someone else’s action? This problem requires a deep understanding of how human poses change over time, learning priors about human appearance and clothing. For example, in Figure 1 the Actor can do a diverse set of actions, from simple actions such as walking and running to more complex actions such as fighting and dancing. For the Imitator, learning a visual prior about their appearance and clothing is essential to animate them at different poses and viewpoints. To tackle this problem, we propose 3DHM, a two-stage framework (see Figure 2) that synthesizes 3D Human Motions by completing a texture map from a single image and then rendering the 3D humans to imitate the actions of the actor.
We use state-of-the-art 3D human pose recovery model 4DHumans [19, 9] for extracting motion signals of the actor, by reconstructing and tracking them over time. Once we have a motion signal in 3D, as a sequence of meshes, one would think we can simply re-texture them with the texture map of the imitator to get an intermediate rendering of the imitation task. However, this requires a complete texture map of the imitator. When given only a single view image of the imitator, we see only a part of their body, perhaps the front side, or the backside but never both sides. To get the complete texture map of the imitator from a single view image, we learn a diffusion model to in-fill the unseen regions of the texture map. This essentially learns a prior about human clothing and appearance. For example, a front-view image of a person wearing a blue shirt would usually have the same color at the back. With this complete texture map, now we can get an intermediate rendering of the imitator doing the actions of the actor. Intermediate rendering means, wrapping the texture map on top of the SMPL [17] mesh to get a body-tight rendering of the imitator.
However, the SMPL [17] mesh renderings are body-tight and do not capture deformations on clothing, like skirts or various hairstyles. To solve this, we learn a second model, that maps from mesh renderings to more realistic images, by controlling the motion with 3D poses. We find out such a simple framework could successfully synthesize realistic and faithful human videos, particularly for long video generations. We show that the 3D control provides a more fine-grained and accurate flow of motion and captures the visual similarities of the imitator faithfully.
While there has been a lot of work on rewriting the motion of an actor [4, 14, 28], each requires either large amounts of data, supervised control signals, or requires careful curations of the training data. For example, Make-a-video [24] can generate decent results while for human videos, it often generates incomplete or nonconsequential videos and fails at faithful reconstruction of humans. Some works [8] use Openpose [6] as intermediate supervision. However, Openpose primarily contains the anatomical key points of humans, it can not be used to indicate the body shape, depth, or other related human body information. DensePose [10] aims to recover highly accurate dense correspondences between images and the body surface to provide dense human pose estimation. However, it can not reflect the texture information from the original inputs. Compared to this line of work, ours fully utilizes the 3D models to control the motion, by providing an accurate dense 3D flow of the motion, and the texture map representation makes it easy to learn appearance prior from a few thousand samples.
2 Related Works
Controllable Human Generation. Human generation is not an easy task. Unlike image translation [16], generating different humans requires the model to understand the 3D structure of the human body. Given arbitrary text prompts or pose conditions [5, 15], we often find out that existing generative models often generate unreasonable human images or videos. Diffusion-HPC [31] proposes a diffusion model with Human Pose Correction and finds that injecting human body structure priors within the generation process could improve the quality of generated images. ControlNet [34] is designed on neural network architecture to control pre-trained large diffusion models to support additional input conditions, such as Openpose [6]. GestureDiffuCLIP [3] designs a neural network to generate co-speech gestures. However, these techniques are not tailored for animating humans, which cannot guarantee the required human appearance and clothing.
Synthesizing Moving People. Synthesizing moving people is very challenging. For example, Make-a-Video [24] or Imagen Video [23] could synthesize videos based on a given instruction. However, the generated video cannot accurately capture human properties correctly and may cause the weird composition of generated humans. Prior methods [8, 29] learn pose-to-pixels mapping directly. However, these designs could only be trained and used for one person. Recent works such as SMPLitex [7] consider human texture estimation from a single image to animate a person. However, there is a visual gap between rendered people via predicted texture map and real humans. Many works start to directly predict pixels based on diffusion models, such as Dreampose [14] and DisCO [28]. DreamPose is controlled by DensePose [10], it aims to synthesize a video containing both human and fabric motion based on a sequence of human body poses. DisCO is directly controlled by Openpose [6], and it aims to animate the human based on the 2D pose information. However, the approach of aligning output pixels for training regularization often leads these models to become overly specialized to certain training data. Moreover, this methodology limits the models’ generalization capabilities, as they often perform well on a few people whose data distribution closely matches that of the training dataset.
Figure 2: Overview of 3DHM: we show an overview of our model pipeline. Given an image of the imitator and a sequence of 3D poses from the actor, we first generate a complete full texture map of the imitator, which can be applied to the 3D pose sequences extracted from the actor to generate texture-mapped intermediate renderings of the imitator. Then we pass these intermediate renderings to the Stage-2 model to project the SMPL mesh rendering to more realistic renderings of real images. Note: red boxes represent inputs, yellow boxes represent intermediate predictions from stage 1, and blue boxes represent the final outputs from stage 2. To create a moving person animation with variable duration and any number of 3D poses, it is only necessary to execute stage 1 once in order to acquire a complete texture map.
3 Synthesizing Moving People
In this section, we discuss our two-stage approach for imitating a motion sequence. Our  3DHM framework embraces the advantage of accurate 3D pose prediction from the state-of-the-art predicting models 4DHumans [19, 9], which could accurately track human motions and extracts 3D human poses of the actor videos. For any given video of the actor we want to imitate, we use 3D reconstruction-based tracking algorithms to extract 3D mesh sequences of the actor. For the inpainting and rendering part, we rely on the pre-trained Stable Diffusion [22] model, which is one of the most recent classes of diffusion models that achieve high competitive results over various generative vision tasks.
Our approach 3DHM is composed of two core parts: Inpainting Diffusion for texture map in-painting as Stage-1 and Rendering Diffusion for human rendering as Stage-2. Figure 2 shows a high-level overview of our framework. In Stage-1, first, for a given single view image, we extract a rough estimate of the texture map by rendering the meshes onto the image and assigning pixels to each visible mesh triangle such that when rendered again it will produce a similar image as the input image. This predicted texture map has only visible parts of the input image. The Stage-1 Diffusion in-painting model takes this partial texture map and generates a complete texture map including the unseen regions. Given this completed texture map, we generate intermediate renderings of SMPL [17] meshes and use Stage-2 model to project the body-tight renderings to more realistic images with clothing. For the Stage-2 Diffusion model, we apply 3D control to animate the imitator to copy the actions of the actor.
3.1 Texture map Inpainting
The goal of Stage-1 model is to produce a plausible complete texture map by inpainting the unseen regions of the imitator. We extract a partially visible texture map by first rendering a 3D mesh onto the input image and sample colors for each visible triangle following 4DHumans [9].
Input. We first utilize a common approach to infer pixel-to-surface correspondences to build an incomplete UV texturemap [32, 7] for texturing 3D meshes from a single RGB image. We also compute a visibility mask to indicate which pixels are visible in 3D and which ones are not.
Target. Since the objective of this modeling is to generate complete texture maps, we generate a pseudo-complete texture map using video data. Since the 4DHumans can track people over time, it continually updates its internal texture map representations as a moving average of visible regions. However to produce more sharp images, for the generative task we found that a median filtering is more suitable than a moving average. While this technique can be applied to any video, in this stage we rely on 2,205 human videos. For each human video, we first extract a partial texture map from each frame. Since each video contains 360 degrees of human views, we calculate a pseudo-complete texture map from a whole video and set it as the target output for Stage 1. In detail, we take the median overall visible parts of texture maps of a video.
Figure 3: Stage-1 of 3DHM: In the first stage, given a single view image of an imitator, we first apply 4Dhumans [9] style sampling approach to extract partial texture map and its corresponding visibility map. These two inputs are passed to the in-painting diffusion model to generate a plausible complete texture map. In this example, while we only see the back view of the imitator, the model was able to hallucinate a plausible front region that is consistent with their clothing.
Model. We finetune directly on the Stable Diffusion Inpainting model [21] that shows great performance on image completion tasks. We input a partial texture map and corresponding visibility mask and obtain the recovered predicted map for the human. We lock the text encoder branch and always feed ‘real human’ as input text of fixed Stable Diffusion models. We refer to our trained model as Inpainting Diffusion. See Figure 3 for the model architecture.
3.2 Human Rendering
In Stage 2, we aim to obtain a realistic rendering of a human imitator doing the actions of the actor. While the intermediate renderings (rendered with the poses from the actor and texture map from Stage-1) can reflect diverse human motion, these SMPL mesh renderings are body-tight and cannot represent realistic rendering with clothing, hairstyles, and body shapes. For example, if we input a scene where a girl is wearing a dress and she is dancing, the intermediate renderings might be able to “dance" but it is impossible to animate the skirt with SMPL mesh rendering. To train this model, in a fully self-supervised fashion, we assume the actor is the imitator, after all a good actor should be a good imitator. This way, we can take any video, and get a sequence of poses from 4DHumans [9] and take any single frame, and get a complete texture map from Stage-1, then get the intermediate renderings by rendering the texture maps on the 3D poses. Now, we have paired data of intermediate renderings and real RGB images. Using this, we collect a large amount of paired data and train our Stage-2 diffusion model with conditioning.
Input: We first apply the generated texture map (fully complete) from Stage 1 to actor 3D body mesh sequences to an intermediate rendering of the imitator performing the actions of the actor. Note at this time, intermediate rendering can only reflect the clothing that fits the 3D mesh (body-tight clothing) but fails to reflect the texture outside the SMPL body, such as the puffed-up region of a skirt, winter jacket, or hat. To obtain the human with complete clothing texture, we input the obtained intermediate renderings and the original image of the person into Rendering Diffusion to render the human in a novel pose with a realistic appearance.
Figure 4: Stage-2 of 3DHM: This figure shows the inference of our Stage-1 approach. Given an intermediate rendering of the imitator with the pose of the actor and the actual RGB image of the imitator, our model can synthesize realistic renderings of the imitator on the pose of the actor.
Target: Since we collected the data by assuming the actor is the imitator, we have the paired data of the intermediate renderings and the real RGB images. This allows us to train this model on lots of data, without requiring any direct 3D supervision.
Model. Similar to ControlNet, we directly clone the weights of the encoder of the Stable Diffusion [20] model as our Controllable branch ("trainable copy") to process 3D conditions. We freeze the pre-trained Stable Diffusion and input noisy latents ($64\times 64$). In the meanwhile, we input a texture mapped 3D human at time $t$ and original human photo input into a fixed VAE encoder and obtain texture mapped 3D human latents ($64\times 64$) and appearance latents ($64\times 64$) as conditioning latents. We feed these two conditioning latents into Rendering Diffusion Controllable branch. The key design principle of this branch is to learn textures from human input and apply them to the texture mapped 3D human during training through the denoising process. The goal is to render a real human with vivid textures from the generated(texture mapped) 3D human from Stage 1. We obtain the output latent and process it to the pixel space via diffusion step procedure and fixed VAE decoder. Same to Stage 1, we lock the text encoder branch and always feed ‘a real human is acting’ as input text of fixed Stable Diffusion models. We refer to our trained model as Rendering Diffusion. In Rendering Diffusion, we predict outputs frame by frame. We show the Stage 2 workflow in Figure 4.
4 Experiments
4.1 Experimental Setup
Dataset. We collect 2,524 3D human videos from 2K2K [11], THuman2.0 [33] and People-Snapshot [2] datasets. 2K2K is a large-scale human dataset with 3D human models reconstructed from 2K resolution images. THuman2.0 contains 500 high-quality human scans captured by a dense DLSR rig. People-Snapshot is a smaller human dataset that captures 24 sequences. We convert the 3D human dataset into videos and extract 3D poses from human videos using 4DHumans. We use 2,205 videos for training and other videos for validation and testing. See the Appendix for more details on the dataset distribution on clothing.
Evaluation Metrics. We evaluate the quality of generated frames of our method with image-based and video-based metrics. For image-based evaluation, we follow the evaluation protocol of DisCO [28] to evaluate the generation quality. We report the average PSNR [13], SSIM [30], FID [12], LPIPS [35], and L1. For video-based evaluation, we use FVD [26]. For pose evaluating 3D pose accuracy we use MPVPE and PA-MVPVE. MPVPE [18], or Mean Per-Vertex Position Error, is a critical metric in 3D human pose estimation, which quantifies the average distance between predicted and actual 3D vertices across a model. This measurement is essential for evaluating the accuracy of 3D reconstructions and pose estimations, with a lower MPVPE indicating higher precision. Complementing this, PA-MPVPE, or Procrustes-Aligned Mean Per-Vertex Position Error, adds another dimension to this evaluation. It involves aligning the predicted and ground truth data using Procrustes Analysis, which neutralizes differences in orientation, scale, and position before calculating the mean error. This alignment allows PA-MPVPE to focus on the structural accuracy of predictions, making it a valuable metric for assessing the relative positioning of vertices in a model, independent of their absolute spatial coordinates.
Implementation Details. As for training all the datasets, we set the constant learning rate as 5e-05 and use the pre-trained diffusion models from diffusers [27] for both Stage-1 and Stage-2. As for Stage 1 Inpainting Diffusion, we finetune on Stable Diffusion Inpainting models [21], which has an 859M total number of trainable parameters and 206M total number of non-trainable parameters, since the VAE is frozen during this stage. We train Rendering Diffusion for 50 epochs and it takes about 2 weeks to run our model on our soup of training datasets. As for Stage 2 Rendering Diffusion, we train the Controllable branch and freeze Stable Diffusion backbones. The total number of trainable parameters in this case is 876M and the total number of non-trainable parameters is 1.1B. We train Rendering Diffusion for 30 epochs and it takes about 2 weeks to run our model on training datasets based on 8 NVIDIA A100 GPUs with a batch size of 4. As for inference, we only need to run Stage-1 once to reconstruct the full texture map of the imitator, and it is used for all other novel poses and viewpoints. We run Stage-2 inference for each frame independently, however since the initial RGB frame of the imitator is conditioned for all frames, the Stage-2 model is able to produce samples that are temporarily consistent.
4.2 Quantitative Results
Baselines. We compare our approaches with past and state-of-the-art methods: DreamPose [14], DisCo [28] and ControlNet [34] (for pose accuracy comparisons) [^1] We utilize the open-source official code and models provided by the authors to implement these baselines. We use diffusers [27] for ControlNet and Openpose extraction, and Detectron2 for DensePose extraction for DisCO. Since  Chan et al. [8] can only work for animating a specific person, we don’t compare with it in this paper.. We set inference steps as 50 for all the approaches for fair comparisons.
Comparisons on Frame-wise Generation Quality. We compare  3DHM with other methods on 2K2K test dataset, which is composed of 50 unseen human videos, at $256\times 256$ resolution. For each human video, we take 30 frames that represent the different viewpoints of each unseen person. The angles range from $0^{\circ}$ to $360^{\circ}$, we take one frame every $12^{\circ}$ to better evaluate the prediction and generalization ability of each model. As for DisCO, we strictly follow their setting and extract OpenPose for inference. As for DreamPose, we extract DensePose for inference. We evaluate the results and calculate the average score over all frames of each video. We set the background as black for all approaches for fair comparisons. We report the average score overall of the same 50 videos and show the comparisons in Table 1. We observe that 3DHM outperforms all the baselines in different metrics.
[…]
DreamPose 35.06 0.80 245.19 0.18 2.12e-04
DisCO 35.38 0.81 164.34 0.15 1.44e-04
Ours 36.18 0.86 154.75 0.12 9.88e-05
Table 1: Quantitative comparison on frame-wise generation quality : We compare our method with prior works on pose condition generation tasks and measure the generation quality of the samples.
Comparisons on Video-level Generation Quality. To verify the temporal consistency of 3DHM, we also report the results following the same test set and baseline implementation as in image-level evaluation. Unlike image-level comparisons, we concatenate every consecutive 16 frames to form a sample of each unseen person on challenging viewpoints. The angles range from $150^{\circ}$ to $195^{\circ}$, we take one frame every $3^{\circ}$ to better evaluate the prediction and generalization ability of each model. We report the average score overall of 50 videos and show the comparisons in Table 2. We observe that 3DHM, though trained and tested by per frame, still embrace significant advantage over prior approaches, indicating superior performance on preserving the temporal consistency with 3D control.
[…]
DreamPose 113.96 950.40
DisCO 83.91 629.18
Ours 55.40 422.38
Table 2: Quantitative comparison on video-level generation quality.
Comparisons on Pose Accuracy. To further evaluate the validity of our model, we estimate 3D poses from generated human videos from different approaches via a state-of-the-art 3D pose estimation model 4DHumans. We use the same dataset setting mentioned above and compare the extracted poses with 3D poses from the target videos. Following the same comparison settings with generation quality, we evaluate the results and calculate the average score over all frames of each video. Beyond DreamPose and DisCO, we also compare with ControlNet, which achieves the state-of-the-art in generating images with conditions, including openpose control. Since ControlNet does not input images, we input the same prompts as ours ‘a real human is acting’ and the corresponding openpose as conditions. We report the average score overall of 50 test videos and show the comparisons in Table 3. We could notice that 3DHM could synthesize moving people following the provided 3D poses with very high accuracy. At the same time, previous approaches might not achieve the same performance by directly predicting the pose-to-pixel mapping. We also notice that 3DHM could achieve superior results on both 2D metrics and 3D metrics, even if DisCO and ControlNet are controlled by Openpose and DreamPose is controlled by DensePose.
Method MPVPE $\downarrow$ PA-MPVPE $\downarrow$
DreamPose 123.07 82.75
DisCO 112.12 63.33
ControlNet 108.32 59.80
Ours 41.08 31.86
Table 3: Quantitative comparison on pose accuracy.
4.3 Qualitative Results
Our work focuses on synthesizing moving people, primarily for clothing and the human body. With the aid of 3D assistance, our approach has the potential to produce human motion videos in various scenarios. We consider challenging 3D poses and motions from 3 sources: 3D human videos, random YouTube videos, and text input.
Poses from Unseen 3D Human Videos. We test our model on different 3D human videos with different human appearances and 3D poses from the 2K2K dataset. We verify that the tested video has never appeared in training data. We display the results in Figure 4(a).
Motions from Random YouTube Videos. We test our model on very different motions from randomly downloaded Youtube videos for an unseen human. We display the results in Figure 4(b).
Motions from Text Inputs. We test our model on motions from arbitrary text prompts. We randomly input an unseen human photo and motions from random text inputs via a widely used human motion generative model (MDM [25]). We display the results in Figure 4(c).
(a) 3DHM with a random human photo and a random 3D pose of various viewpoints. We show that even if the person’s photo is from a side angle, our stage 1 can help reconstruct the full texture map, which could be used to obtain full body information. Stage 2 can add texture information based on a given input.
(b) 3DHM with a random human photo and motions from random YouTube Videos. This example is from Gene Kelly’s dancing video.
(c) 3DHM with a random human photo and motions generated from text inputs by MDM, a Human Motion Diffusion Model [25].
Figure 5: Qualitative results on different viewpoints of the same pose; motions from random videos and motions from text input.
5 Analysis and Discussion
5.1 Ablation Study
To further verify the components of our methods, we train on training dataset and test on test datasets. We extract the 3D rendered pose from these 50 test video tracks. Same with the settings in quantitative comparison, we calculate the average scores of PSNR, SSIM, VGG, L1, LPIPS among all the generated frames and targeted original frames and report the results on both frame-wise metric (Table 4), video-level metric (Table 5) and pose accuracy (Table 6). We find that both texture map reconstruction and appearance latents are critical to the model performance. Also, we notice that directly adding SMPL parameters into the model during training may not bring improved performance considering all evaluation metrics. This is presumably due to the imprecision of SMPL parameters, which could provide contradictory information throughout the diffusion training process if they are not incorporated correctly.
[…]
Default 36.18 0.86 154.75 0.12 9.88e-05
w/o Texture map 35.00 0.78 237.42 0.20 2.35e-04
w/o Appearance Latents 36.07 0.86 167.58 0.12 1.03e-04
adding SMPL parameters 36.42 0.87 157.60 0.12 8.87e-05
Table 4: Ablation study of Rendering Diffusion. We compare the frame-wise generation quality under different settings. We notice both texturemap reconstruction and appearance latents are critical to the model performance.
[…]
Default 55.40 422.38
w/o Texture map 113.97 632.67
w/o Appearance Latents 93.21 715.51
adding SMPL parameters 72.35 579.90
Table 5: Ablation study of Rendering Diffusion. We compare the video-level generation quality under different settings. We notice that although adding SMPL parameters achieve better performance on frame-wise setting but may yield worse temporal consistency than default settings.
Method MPVPE $\downarrow$ PA-MPVPE $\downarrow$
Default 41.08 31.86
w/o Texture map 92.94 59.18
w/o Appearance Latents 41.99 32.82
adding SMPL parameters 39.16 29.67
Table 6: Ablation study of Rendering Diffusion. We compare the pose accuracy under different settings.
5.2 2D Control and 3D Control
We also compare the results of the official model from DreamPose and DisCO on a random person on a random real human photo which ensures distinct data distribution. We display the qualitative results of various viewpoints in Figure 6. DreamPose, DisCO, and 3DHM are all initialize the U-Net model with the pre-trained weights of Stable Diffusion. We notice that 3DHM can generalize well to unseen real humans though it is only trained by limited 3D humans. Since DreamPose requires subject-specific finetuning of the UNet to achieve better results, it cannot directly generalize well on a random human photo. As for DisCO, though it has been trained with an effective human attribute pre-training on multiple public datasets for better generalizability to unseen humans, still fails to synthesize people without the target pose. We assume this is because 3DHM adds rigid 3D control to better correlate the appearance to the poses, and preserve the body shape. Training with OpenPose or DensePose cannot guarantee the mapping between textures and poses, which makes it hard for the models to generalize.
Figure 6: Qualitative comparison with other 2D control approaches on a random real human photo (a Korean actress). We apply various 3D poses or the same 3D pose from different viewpoints. It could be noticed that 2D poses may not be able to capture the folding motion, and details of the human body. We could notice that our approach 3DHM could bridge this gap with 3D control.
5.3 Limitations
As 3DHM generates the frames of the human motion videos independently, there is no guarantee of consistency in time. For example, the clothing light may change between consecutive frames. One possible solution is to train the model to predict multiple frames simultaneously. Another possible solution is to condition the generation process on previously generated frames via stochastic conditioning [1]. Additionally, since 3DHM is trained on a dataset of 2K people, not all the detailed textures can be reconstructed completely during inference (e.g. unique logos on the clothes). We hypothesize this could be alleviated by training with more human data.
6 Conclusion
In this paper, we propose 3DHM, a two-stage diffusion model-based framework that enables synthesizing moving people based on one random photo and target human poses. A notable aspect of our approach is that we employ a cutting-edge 3D pose estimation model to generate human motion data, allowing our model to be trained on arbitrary videos without necessitating ground truth labels. Our method is suitable for long-range motion generation, and can deal with arbitrary poses with superior performance over previous approaches.
Acknowledgement
We thank the Machine Common Sense project and ONR MURI award number N00014-21-1-2801. We also thank Google’s TPU Research Cloud (TRC) for providing cloud TPUs. We thank Georgios Pavlakos, Shubham Goel, and Jane Wu for the constructive feedback and helpful discussions.
Appendix A Dataset Analysis
Figures 6(a) and 6(b) present the clothing type statistics of the training data (2,205 humans) and test data (50 humans). We count people based on four clothing categories: skirted attire, suit, casual wear, and others. In some cases, the clothing belongs to skirted attire and suits or casual wear, we will count this as skirted attire. For each clothing category, we tally two styles: tight-fitting and loose-fitting.
In this paper, we only train on limited human videos, we assume training with more human videos could largely boost the model generalization on the fly. Given that 3DHM makes use of a cutting-edge 3D pose estimation model and only requires human videos without additional labels for training, it could be trained with numerous and any human videos such as movies, etc.
Appendix B 3DHM Training Features
As has been mentioned in the paper, 3DHM is in a fully self-supervised fashion. Here we summarize the key training features of our approach:
  • 3DHM training pipeline (for both stages) is self-supervised.
  • 3DHM does not use any additional annotations. It is trained with pseudo-ground-truth as we use cutting-edge software which can detect, segment, track and 3Dfy humans (H4D).
  • 3DHM is scalable and its scaling can be done readily in the future given additional videos of humans in motion and computing resources.
(a) Training data distribution.
(b) Testing data distribution.
Figure 7: Data distribution. We split the clothing type into 4 categories: skirted attire, suit, casual wear, and others. We split each category into two types: loose and tight. We report the number of each category and type and display the overall distribution. We could notice that most clothing is casual wear and a large portion belongs to tight-fitting.