### v2

# Synthesizing Moving People with 3D Control

Abstract
In this paper, we present a diffusion model-based framework for animating people from a single image for a given target 3D motion sequence. Our approach has two core components: a) learning priors about invisible parts of the human body and clothing, and b) rendering novel body poses with proper clothing and texture. For the first part, we learn an in-filling diffusion model to hallucinate unseen parts of a person given a single image. We train this model on texture map space, which makes it more sample-efficient since it is invariant to pose and viewpoint. Second, we develop a diffusion-based rendering pipeline, which is controlled by 3D human poses. This produces realistic renderings of novel poses of the person, including clothing, hair, and plausible in-filling of unseen regions. This disentangled approach allows our method to generate a sequence of images that are faithful to the target motion in the 3D pose and, to the input image in terms of visual similarity. In addition to that, the 3D control allows various synthetic camera trajectories to render a person. Our experiments show that our method is resilient in generating prolonged motions and varied challenging and complex poses compared to prior methods. Please check our website for more details: 3DHM.github.io.
Figure 1: The Imitation Game: Given a video of a person "The Actor", we want to transfer their motion to a new person "The Imitator". In this figure, the first row shows a sequence of frames of the actor from a ballerina Dance of the Sugar Plum Fairy. The inset row shows the 3D poses extracted from this video. Now, given any single image of a new person The Imitator, our model can synthesize new renderings of the imitator, to copy the actions of the actor in 3D.
1 Introduction
Given a random photo of a person, can we accurately animate that person to imitate someone else’s action? This problem requires a deep understanding of how human poses change over time, learning priors about human appearance and clothing. For example, in Figure 1 the Actor can do a diverse set of actions, from simple actions such as walking and running to more complex actions such as fighting and dancing. For the Imitator, learning a visual prior about their appearance and clothing is essential to animate them at different poses and viewpoints. To tackle this problem, we propose 3DHM, a diffusion framework (see Figure 2) that synthesizes 3D Human Motions by completing a texture map from a single image and then rendering the 3D humans to imitate the actions of the actor.
Figure 2: Overview of 3DHM: we show an overview of our model pipeline. Given an image of the imitator and a sequence of 3D poses from the actor, we first generate a complete full texture map of the imitator, which can be applied to the 3D pose sequences extracted from the actor to generate texture-mapped intermediate renderings of the imitator. Then we pass these intermediate renderings to the Stage-2 model to project the SMPL mesh rendering to more realistic renderings of real images.
We use state-of-the-art 3D human pose recovery model 4DHumans [24, 9] for extracting motion signals of the actor, by reconstructing and tracking them over time. Once we have a motion signal in 3D, as a sequence of meshes, one would think we can simply re-texture them with the texture map of the imitator to get an intermediate rendering of the imitation task. However, this requires a complete texture map of the imitator. When given only a single view image of the imitator, we see only a part of their body, perhaps the front side, or the backside but never both sides. To get the complete texture map of the imitator from a single view image, we learn a diffusion model to in-fill the unseen regions of the texture map. This essentially learns a prior about human clothing and appearance. For example, a front-view image of a person wearing a blue shirt would usually have the same color at the back. With this complete texture map, now we can get an intermediate rendering of the imitator doing the actions of the actor. Intermediate rendering means, wrapping the texture map on top of the SMPL [20] mesh to get a body-tight rendering of the imitator.
However, the SMPL [20] mesh renderings are body-tight and do not capture deformations on clothing, like skirts or various hairstyles. To solve this, we learn a second model, that maps from mesh renderings to more realistic images, by controlling the motion with 3D poses. We find out such a simple framework could successfully synthesize realistic and faithful human videos, particularly for long video generations. We show that the 3D control provides a more fine-grained and accurate flow of motion and captures the visual similarities of the imitator faithfully.
While there has been a lot of work on rewriting the motion of an actor [3, 16, 32], each requires either large amounts of data, supervised control signals, or requires careful curations of the training data. For example, Make-a-video [29] can generate decent results while for human videos, it often generates incomplete or nonconsequential videos and fails at faithful reconstruction of humans. Some works [7] use Openpose [5] as intermediate supervision. However, Openpose primarily contains the anatomical key points of humans, it can not be used to indicate the body shape, depth, or other related human body information. DensePose [10] aims to recover highly accurate dense correspondences between images and the body surface to provide dense human pose estimation. However, it can not reflect the texture information from the original inputs. Compared to this line of work, ours fully utilizes the 3D models to control the motion, by providing an accurate dense 3D flow of the motion, and the texture map representation makes it easy to learn appearance prior from a few thousand samples.
2 Related Works
Controllable Human Generation. Human generation is not an easy task. Unlike image translation [18], generating different humans requires the model to understand the 3D structure of the human body. Given arbitrary text prompts or pose conditions [4, 17], we often find out that existing generative models often generate unreasonable human images or videos. Diffusion-HPC [36] proposes a diffusion model with Human Pose Correction and finds that injecting human body structure priors within the generation process could improve the quality of generated images. ControlNet [41] is designed on neural network architecture to control pre-trained large diffusion models to support additional input conditions, such as Openpose [5]. GestureDiffuCLIP [2] designs a neural network to generate co-speech gestures. However, these techniques are not tailored for animating humans, and do not guarantee the required human appearance and clothing.
Synthesizing Moving People. Synthesizing moving people is very challenging. For example, Make-a-Video [29] or Imagen Video [28] could synthesize videos based on a given instruction. However, the generated video cannot accurately capture human properties correctly and may cause the weird composition of generated humans. Older methods [7, 33] learn pose-to-pixels mapping directly, but they require being trained for each new person separately. Recent works such as SMPLitex [6] consider human texture estimation from a single image to animate a person. However, there is a visual gap between rendered people via predicted texture map and real humans. Many works start to directly predict pixels based on diffusion models, such as Dreampose [16], DisCO [32], AnimateAnyone [15], MagicAnimate [38], Champ [43], etc. [8, 34, 21, 23]. DreamPose and MagicAnimate is controlled by DensePose [10], it aims to synthesize a video containing both human and fabric motion based on a sequence of human body UV or Segmentation maps. DisCO and AnimateAnyone is directly controlled by Openpose [5], and it aims to animate the human based on the 2D pose information. Champ [43] utilizes the multiple condition maps rendered from SMPL mesh to further enhance detailed controllability. However, the approach of aligning output pixels for training regularization often leads these models to become overly specialized to certain training data. Moreover, this methodology limits the models’ generalization, as they often perform well on a few people whose data distribution closely matches that of the training dataset.
3 Synthesizing Moving People
In this section, we discuss our two-stage approach for imitating a motion sequence. Our  3DHM framework embraces the advantage of accurate 3D pose prediction from the state-of-the-art predicting models 4DHumans [24, 9], which could accurately track human motions and extracts 3D human poses of the actor videos. For any given video of the actor we want to imitate, we use 3D reconstruction-based tracking algorithms to extract 3D mesh sequences of the actor. For the inpainting and rendering part, we rely on the pre-trained Stable Diffusion [27] model, which is one of the most recent classes of diffusion models that achieve high competitive results over various generative vision tasks.
Our approach 3DHM is composed of two core parts: Inpainting Diffusion for texture map in-painting as Stage-1 and Rendering Diffusion for human rendering as Stage-2. Figure 2 shows a high-level overview of our framework. In Stage-1, first, for a given single view image, we extract a rough estimate of the texture map by rendering the meshes onto the image and assigning pixels to each visible mesh triangle such that when rendered again it will produce a similar image as the input image. This predicted texture map has only visible parts of the input image. The Stage-1 Diffusion in-painting model takes this partial texture map and generates a complete texture map including the unseen regions. Given this complete texture map, we generate intermediate renderings of SMPL [20] meshes and use Stage-2 model to project the body-tight renderings to more realistic images with clothing. For the Stage-2 model, we apply 3D control to animate the imitator to copy the actions of the actor.
3.1 Texture map Inpainting
The goal of Stage-1 model is to produce a plausible complete texture map by inpainting the unseen regions of the imitator. We extract a partially visible texture map by first rendering a 3D mesh onto the input image and sample colors for each visible triangle following 4DHumans [9].
Input. We first utilize a common approach to infer pixel-to-surface correspondences to build an incomplete UV texturemap [37, 6] for texturing 3D meshes from a single RGB image. We also compute a visibility mask to indicate which pixels are visible in 3D and which ones are not.
Target. We train our model on a large 3D human texture dataset [19], which contains 50k high-fidelity textured UV map of SMPL [20]. To strengthen the model’s 3D geometry consistency in completing the partial texturemap, We densely sample a group of visibility masks from 360 degrees of SMPL mesh, which then mask out Ground-Truth texture map to produce the pseudo-partial texture map during training the inpainting model. Benefiting from the extensive collection of texture maps from diverse human appearances, as well as the numerous visibility masks from various viewpoints.
Figure 3: Stage-1 of 3DHM: In the first stage, given a single view image of an imitator, we first apply 4Dhumans [9] style sampling approach to extract partial texture map and its corresponding visibility map. These two inputs are passed to the in-painting diffusion model to generate a plausible complete texture map. In this example, while we only see the front view of the imitator, the model was able to hallucinate a plausible back region that is consistent with their clothing.
Model. We finetune directly on the Stable Diffusion Inpainting model [26] that shows great performance on image completion tasks. Given a single RGB human image, we predict the human mesh and calculate its corresponding visibility mask and partial texture map, which is then recovered by the in-painting model to complete texture map for the human. We lock the text encoder branch during training and feed "3D realistic human, UV texturemap" as input text condition. We refer to our trained model as Inpainting Diffusion. See Figure 3 for the model architecture.
3.2 Human Rendering
In Stage 2, we aim to obtain a realistic rendering of a human imitator doing the actions of the actor. While the intermediate renderings (rendered with the poses from the actor and texture map from Stage-1) can reflect diverse human motion, these SMPL mesh renderings are body-tight and cannot represent realistic rendering with clothing, hairstyles, and body shapes. We train a model for realistic rendering, in a fully self-supervised fashion, by relying on the actor as the imitator. We obtain a sequence of poses from 4DHumans [9] for each training video and use Stage-1 on single frames to obtain a complete texture map. We then pair the intermediate renderings (i.e. the rendered texture maps on the 3D poses) with the original frames from which they were obtained. We collect a large amount of paired data and train our Stage-2 diffusion model with conditioning.
Input: We first apply the generated complete texture map from Stage-1 to the actor’s 3D body mesh sequences to obtain the intermediate rendering. Note that the rendering can only reflect the clothing that fits the 3D mesh (body-tight clothing) but fails to reflect the texture outside the SMPL body (e.g., the puffed-up skirt region, or hat). To obtain the human with complete clothing texture, we input the obtained intermediate renderings and the original image of the person into Rendering Diffusion to render the human in a novel pose with a realistic appearance.
Target: Since we collected the data by assuming the actor is the imitator, we have the paired data of the intermediate renderings and the real RGB images. This allows us to train this model on lots of data, without requiring any direct 3D supervision.
Model. Similar to ControlNet, we directly clone the weights of the encoder of the Stable Diffusion [25] model as our Controllable branch ("trainable copy") to process 3D conditions. We freeze the pre-trained Stable Diffusion. In the meanwhile, we input a texture-mapped 3D human at time $t$ and original human photo input into a fixed VAE encoder and obtain texture-mapped 3D human latents ($64\times 64$) and appearance latents ($64\times 64$) as conditioning latents. We feed these two conditioning latents into Rendering Diffusion Controllable branch. The key design principle of this branch is to learn textures from human input and apply them to the texture-mapped 3D human during training through the denoising process. The goal is to render a real human with vivid textures from the generated(texture-mapped) 3D human from Stage 1. We obtain the output latent and process it to the pixel space via diffusion step procedure and fixed VAE decoder. We refer to our trained model as Rendering Diffusion. In Rendering Diffusion, we predict outputs frame by frame. We show the Stage 2 workflow in Figure 4.
Figure 4: Stage-2 of 3DHM: Given an intermediate rendering of the imitator with the pose of the actor and the actual RGB image of the imitator, our model can synthesize realistic renderings of the imitator on the pose of the actor.
4 A Complete Approach
In this section, we discuss how to scale up our method to real-world domains. We first discuss the challenges of visual appearance injection for Diffusion models and then propose a novel framework to enhance consistent appearance alignment. For any given reference human images (the imitators), our network can generate high-fidelity results benefiting from the enhanced appearance encoder [39]. To ensure the visual consistency of both human identity and background from reference images, we use a trainable copy of base Stable Diffusion model [27] to inject appearance information that perfectly aligns with the backbone denoising Diffusion model. To encourage temporally smooth reconstructions, we utilize a temporal diffusion model [11] to learn the temporal correlation within motion sequences. To generate smooth and consistent videos, we propose an easy but efficient method that takes the previous frame for consequential video generation. The detailed framework is shown in Figure 5.
4.1 Enhance Appearance Alignment
The key challenge in scaling up our method to real-world domains is to both maintain the complex background and the human appearance from reference images consistently. The Stable Diffusion Model [27] is trained on text-to-image tasks, and focuses on semantics of the generated image. However, in our Stage-2 rendering task, instead of semantic features, the model needs more low-level visual features to reconstruct the detailed structure and appearance from the input images. Given the imitator’s images as reference, our approach simultaneously leverages the capabilities of Stable Diffusion and uses the reference image prompts to obtain more accurate generation. We use a lightweight image adapter [39] to condition diffusion on image prompts. We also add a trainable branch of the Stable Diffusion model, namely Reference Net, to enhance appearance consistency on both the input image’s background and human appearance.
Input. Same with Stage-2, we input the generated imitator’s texture map from Stage-1 with actor’s 3D mesh sequences to get the intermediate rendering. Then the intermediate rendering is input to the 3D Controllable branch as motion condition. The original imitator’s RGB image is input to the Reference Net and the image adaptor as the appearance guidance.
Target. We scale up our model on about 1000 real human videos collected from the Internet, each containing 2-10 seconds solo dancing. For this stage, we keep the assumption that: imitator and actor are identical and randomly sample the imitator and actor. The objective now is to drive the reference imitator image with actor’s pose to generate the corresponding target image. We trained on our real and virtual datasets together to teach model focus on complex pose and appearance variance, and the 3D view consistency respectively.
Figure 5: Scaled up Stage-2 of 3DHM Model: To enable consistent background and human generation, we train ReferenceNet with ControlNet, and then only finetune the temporal-attention layer of the UNet and keep other components frozen.
Model. To better align the input image’s appearance with the denoising backbone, we make a trainable copy of the pre-trained Stable Diffusion as our ReferenceNet. Notice that different from the Section 3.2, instead of adding appearance latents with 3D human latents together, we now separately input the imitator’s appearance latents to ReferenceNet to extract level of details appearance latents. The appearance latents are then injected into the denoising backbone to condition the denoising process, which help to keep consistent background and human appearance for different poses. The pretrained IP-adapter [39] is also integrated into all branches to keep human identity.
4.2 Temporal Consistency
Once our image-to-image model learned to generate the imitator frames we can apply it frame-by-frame for image-to-video generation. However, the generated video may suffer from jittering due to the lack of temporal consistency. Based on the Stage-2 model, we adapt a temporal model pretrained on a large video dataset  [11] to learn the motion’s coherence and appearance consistency. Though previous works  [32, 16, 15] also have similar temporal layers and success for short video, they still suffer from inconsistency in long video generation results. To effectively solve the unstable and randomness between each short video clip generated from temporal sliding windows, we design an easy process to concatenate the previous frame’s latents with the consequential video.
Input. In this stage, our model is optimized directly on video data. During training, for each video clip, we extract $F$ consecutive frames as the target of actor’s motion sequence and randomly pick a frame as the imitator’s reference image. Now the 3D Control branch takes $F$ consecutive intermediate rendering as motion sequences to drive the imitator’s image to generate a temporal coherent video. Notice that our Reference Net here will not cost extra computing time since there is only one reference image for each video clip.
Model. The short video clip is now defined as $V\in\mathbb{R}^{B\times C\times F\times H\times W}$, with batch size $B$, the number of channels $C$, the number of consecutive frames $F$, height $H$ and width $W$ respectively. The temporal layers are inserted at each resolution level. For level $i$ , the 5D latents $v_{i}\in\mathbb{R}^{b\times c\times f\times h\times w}$ is reshaped to ${(b\times h\times w)\times f\times c}$ within the temporal layer for self-attention to align feature maps across frames. The temporal attention mechanism above not only effectively smooths the flickering and jittering, but also improves the motion and appearance consistency in generated videos. However, during long video generation, since the video clips are generated independently and concatenated together, the different random noises will cause inconsistency between each video clip. To facilitate the cross-clips consistency, we take the last frame $V_{k}^{f}$ from $k$-th generated video clip $V_{k}^{1:f}$ to condition on the next video clip generation $V_{k+1}^{1:f}$. The $V_{k}^{f}$ is input to the Reference Net to extract corresponding latents for each resolution level $i$, and then fed into the temporal layers concatenated with original 5D latents along the temporal dimension. The conditioned temporal layers at level $i$ now attention across latent $v_{i}^{0:f}\in\mathbb{R}^{b\times c\times(1+f)\times h\times w}$ and then trunck the previous frame $v_{i}^{0}$ to get the conditioned results $\bar{v}_{i}^{1:f}$. During this stage, we freeze all other parameters and train only the temporal model. Zero-initialize [41] is also applied to the temporal layers to eliminate harmful noise during training.
5 Experiments
5.1 Experimental Setup
Dataset. We collect 2,524 3D human videos from 2K2K [12], THuman2.0 [40] and People-Snapshot [1] datasets. 2K2K is a large-scale human dataset with 3D human models reconstructed from 2K resolution images. THuman2.0 contains 500 high-quality human scans captured by a dense DLSR rig. People-Snapshot is a smaller human dataset that captures 24 sequences. We convert the 3D human dataset into videos and extract 3D poses from human videos using 4DHumans. We use 2,205 videos for training and other videos for validation and testing. See the Appendix for more details on the dataset distribution on clothing.
Evaluation Metrics. We evaluate the quality of generated frames of our method with image-based and video-based metrics. For image-based evaluation, we follow the evaluation protocol of DisCO [32] to evaluate the generation quality. We report the average PSNR [14], SSIM [35], FID [13], LPIPS [42], and L1. For video-based evaluation, we use FVD [30]. For pose evaluating 3D pose accuracy we use Mean Per-Vertex Position Error (MPVPE) and Procrustes-Aligned Mean Per-Vertex Position Error (PA-MVPVE [22]).
Implementation Details. We set a learning rate of 5e-05 and use the pre-trained diffusion models for both stages. For Stage 1 Inpainting Diffusion, we fine-tune Stable Diffusion Inpainting models [26] We train Rendering Diffusion for 50 epochs (requires about 2 weeks on our compute). For Stage 2 Rendering Diffusion, we train the Controllable branch and freeze Stable Diffusion backbones. The total number of trainable parameters in this case is 876M. We train Rendering Diffusion for 30 epochs (requires about 2 weeks on 8 NVIDIA A100 GPUs with a batch size of 4). During inference, we only need to run Stage-1 once to reconstruct the full texture map of the imitator, which is used for all other novel poses and viewpoints. We run Stage-2 inference for each frame independently, however since the initial RGB frame of the imitator is conditioned for all frames, the Stage-2 model is able to produce samples that are temporarily consistent.
(a) 3DHM with random 3D poses from various viewpoints. We show that even if the person’s photo is from a side angle, our stage 1 can help reconstruct the full texture map, which could be used to obtain full body information. Stage 2 can add texture information based on a given input.
(b) 3DHM with motions from random YouTube Videos.
Figure 6: Qualitative results on different viewpoints of the same pose and motions from random videos.
5.2 Quantitative Results
Baselines. We compare our approaches with past and state-of-the-art methods: DreamPose [16], DisCo [32] and ControlNet [41] (for pose accuracy comparisons) [^1] We utilize the open-source official code and models provided by the authors to implement these baselines. We use diffusers [31] for ControlNet and Openpose extraction, and Detectron2 for DensePose extraction for MagicAnimate and DreamPose. Since  Chan et al. [7] can only work for animating a specific person, we don’t compare with it in this paper.. We set inference steps as 50 for all the approaches for fair comparisons.
Comparisons on Frame-wise Generation Quality. We compare  3DHM with other methods on 2K2K test dataset, which is composed of 50 unseen human videos, at $256\times 256$ resolution. For each human video, we take 30 frames that represent the different viewpoints of each unseen person. The angles range from $0^{\circ}$ to $360^{\circ}$, we take one frame every $12^{\circ}$ to better evaluate the prediction and generalization ability of each model. As for DisCO, we strictly follow their setting and extract OpenPose for inference. We extract DensePose for inference DreamPose and MagicAnimate. We evaluate the results and calculate the average score over all frames of each video. We set the background as black for all approaches for fair comparisons. We report the average score of the same 50 videos and show the comparisons in Table 1. We observe that 3DHM outperforms all the baselines in different metrics.
[…]
DreamPose 35.06 0.80 245.19 0.18 2.12e-04 113.96 950.40
DisCO 35.38 0.81 164.34 0.15 1.44e-04 83.91 629.18
MagicAnimate 32.57 0.65 300.66 0.29 5.80E-04 140.45 900.70
Ours 36.18 0.86 154.75 0.12 9.88e-05 55.40 422.38
Table 1: Quantitative comparison on generation quality: We compare our method with prior works on pose condition generation tasks and measure the generation quality of the samples.
Comparisons on Video-level Generation Quality. To verify the temporal consistency of 3DHM, we also report the results following the same test set and baseline implementation as in image-level evaluation. Unlike image-level comparisons, we concatenate every consecutive 16 frames to form a sample of each unseen person on challenging viewpoints. The angles range from $150^{\circ}$ to $195^{\circ}$, we take one frame every $3^{\circ}$ to better evaluate the prediction and generalization ability of each model. We report the average score overall of 50 videos and show the comparisons in Table 1. We observe that 3DHM, though trained and tested by per frame, still embraces significant advantage over prior approaches, indicating superior performance on preserving the temporal consistency with 3D control.
Comparisons on Pose Accuracy. To further evaluate the validity of our model, we estimate 3D poses from generated human videos from different approaches via a state-of-the-art 3D pose estimation model 4DHumans. We use the same dataset setting mentioned above and compare the extracted poses with 3D poses from the target videos. Following the same comparison settings with generation quality, we evaluate the results and calculate the average score over all frames of each video. Beyond DreamPose and DisCO, we also compare with ControlNet, which achieves the state-of-the-art in generating images with conditions, including openpose control. Since ControlNet does not input images, we input the same prompts as ours ‘a real human is acting’ and the corresponding openpose as conditions. We report the average score overall of 50 test videos and show the comparisons in Table 2. We could notice that 3DHM could synthesize moving people following the provided 3D poses with very high accuracy. At the same time, previous approaches might not achieve the same performance by directly predicting the pose-to-pixel mapping. We also notice that 3DHM could achieve better results on both 2D metrics and 3D metrics, even if DisCO and ControlNet are controlled by Openpose and DreamPose is controlled by DensePose.
Method MPVPE $\downarrow$ PA-MPVPE $\downarrow$
DreamPose 123.07 82.75
DisCO 112.12 63.33
ControlNet 108.32 59.80
Ours 41.08 31.86
Table 2: Quantitative comparison on pose accuracy. We measure the pose accuracy in the generated images given the conditioned pose as the ground truth. We can see that our model is very accurate in persevering the poses in generated images.
5.3 Ablation Study
To further verify the components of our methods, we train on training dataset and test on test datasets. We extract the 3D rendered pose from these 50 test video tracks. Same with the settings in quantitative comparison, we calculate the average scores among all the generated frames and targeted original frames and report the results on both frame-wise metric (PSNR, SSIM, FID, LPIPS, L1), video-level metric (FID-VID, FVD) and pose accuracy (MPVPE, PA-MPVPE) in Table 4. We find that both texture map reconstruction and appearance latents are critical to the model performance. Also, we notice that directly adding SMPL parameters into the model during training may not bring improved performance considering all evaluation metrics.
Running Cost. Here we outline the comparison of parameters and running time with other methods in Table 1 using a single GPU A100. We show the comparison in Table 3.
Method Time (second/frame) Parameter
DreamPose 22.0 1.0B
DisCO 5.0 2.0B
MagicAnimate 10.0 2.0B
Ours 3.2 2.0B
Table 3: Comparison of running cost. We compare inference time for different models, and we can see that our models is faster in comparison with our models.
6 Analysis and Discussion
[…]
Default 36.18 0.86 154.75 0.12 9.88e-05 55.40 422.38 41.08 31.86
w/o Texture map 35.00 0.78 237.42 0.20 2.35e-04 113.97 632.67 92.94 59.18
w/o Appearance Latents 36.07 0.86 167.58 0.12 1.03e-04 93.21 715.51 41.99 32.82
adding SMPL parameters 36.42 0.87 157.60 0.12 8.87e-05 72.35 579.90 39.16 29.67
Table 4: Ablation study of Rendering Diffusion. We compare the frame-wise generation quality, video-level generation quality and the pose accuracy under different settings. We notice both texturemap reconstruction and appearance latents are critical to the model performance. The results show that although adding SMPL parameters achieve better performance on frame-wise setting but may yield worse temporal consistency than default settings. Note: we use bold to represent the best result and underline to represent the second-best result.
6.1 Qualitative Results
Our work focuses on synthesizing moving people, primarily for clothing and the human body. With the aid of 3D assistance, our approach has the potential to produce human motion videos in various scenarios. We consider challenging 3D poses and motions from 2 sources: 3D human videos and random YouTube videos. We utilize our model, which has been scaled up for real-world domains.
Poses from Unseen 3D Human Videos. We test our model on different 3D human videos with different human appearances and 3D poses from the 2K2K dataset. We verify that the tested video has never appeared in training data. We display the results in Figure 6(a).
Motions from Random YouTube Videos. We test our model on very different motions from randomly downloaded YouTube videos for an unseen human. We display the results in Figure 6(b). The results show that 3DHM can efficiently animate any person using random motion resources, accurately following the 3D poses from challenging motion sources.
6.2 Qualitative Comparison
We also compare the results of the official model from DreamPose, DisCO, and MagicAnimate on a random person on a random real human photo which ensures distinct data distribution. We display the qualitative results of various poses on real human photos in Figure 7. We notice that 3DHM can generalize well to unseen real humans though it is only trained by limited 3D humans. Since DreamPose requires subject-specific finetuning of the UNet to achieve better results, it cannot directly generalize well on a random human photo. As for DisCO, though it has been trained with an effective human attribute pre-training on multiple public datasets for better generalizability to unseen humans, still fails to synthesize people without the target pose. MagicAnimate uses 3D pose features (DensePose) which better controls the appearance of input images. However it always suffers from severe artifacts on DensePose segmentation maps, which severely ruins the pose accuracy and consistency. We assume this is because 3DHM adds rigid 3D control to better correlate the appearance to the poses, and preserve the body shape. Training with OpenPose or DensePose cannot guarantee the mapping between textures and poses, which makes it hard for the models to generalize.
Figure 7: Qualitative comparison with other state-of-the-art approaches on a real human photo.
6.3 Limitations
As 3DHM has been trained with limited data (around 2K synthetic humans and 1K real humans), it might struggle to predict the texture details of the unseen side of the input human photo. However, we believe this issue can be mitigated by scaling up with more human data.
7 Conclusion
In this paper, we propose 3DHM, a two-stage diffusion model-based framework that enables synthesizing moving people based on one random photo and target sequence of human poses. A notable aspect of our approach is that we employ a cutting-edge 3D pose estimation model to generate human motion data, allowing our model to be trained on arbitrary videos without necessitating ground truth labels. Our method is suitable for long-range motion generation, and can deal with arbitrary poses with superior performance over previous approaches, by preserving the poses of the target motion, and clothing, face identities and smoother motion between frames.
Acknowledgement
We thank the Machine Common Sense project and ONR MURI award number N00014-21-1-2801. We also thank Google’s TPU Research Cloud (TRC) for providing cloud TPUs. We thank Georgios Pavlakos, Shubham Goel, Jane Wu and Karttikeya Mangalam for the constructive feedback and helpful discussions.
Appendix A Dataset Analysis
Figures 8(a) and 8(b) present the clothing type statistics of the synthetic training data (2,205 humans) and test data (50 humans). We count people based on four clothing categories: skirted attire, suit, casual wear, and others. In some cases, the clothing belongs to skirted attire and suits or casual wear, we will count this as skirted attire. For each clothing category, we tally two styles: tight-fitting and loose-fitting.
In this paper, we only train on limited human videos, we assume training with more human videos could largely boost the model generalization on the fly. Given that 3DHM makes use of a cutting-edge 3D pose estimation model and only requires human videos without additional labels for training, it could be trained with numerous and any human videos such as movies, etc.
Appendix B 3DHM Training Features
As has been mentioned in the paper, 3DHM is in a fully self-supervised fashion. Here we summarize the key training features of our approach:
  • 3DHM training pipeline (for both stages) is self-supervised.
  • 3DHM does not use any additional annotations. It is trained with pseudo-ground-truth as we use cutting-edge software which can detect, segment, track and 3Dfy humans (H4D).
  • 3DHM is scalable and its scaling can be done readily in the future given additional videos of humans in motion and computing resources.
(a) Training data distribution.
(b) Testing data distribution.
Figure 8: Data distribution. We split the clothing type into 4 categories: skirted attire, suit, casual wear, and others. We split each category into two types: loose and tight. We report the number of each category and type and display the overall distribution. We could notice that most clothing is casual wear and a large portion belongs to tight-fitting.